# Bipedal Safe Navigation over Uncertain Rough Terrain: Unifying Terrain Mapping and Locomotion Stability

### v1

@@ Abstract
We study the problem of bipedal robot navigation in complex environments with uncertain and rough terrain. In particular, we consider a scenario in which the robot is expected to reach a desired goal location by traversing an environment with uncertain terrain elevation. Such terrain uncertainties induce not only untraversable regions but also robot motion perturbations. Thus, the problems of terrain mapping and locomotion stability are intertwined. We evaluate three different kernels for Gaussian process (GP) regression to learn the terrain elevation. We also learn the motion deviation resulting from both the terrain as well as the discrepancy between the reduced-order Prismatic Inverted Pendulum Model used for planning and the full-order locomotion dynamics. We propose a hierarchical locomotion-dynamics-aware sampling-based navigation planner. The global navigation planner plans a series of local waypoints to reach the desired goal locations while respecting locomotion stability constraints. Then, a local navigation planner is used to generate a sequence of dynamically feasible footsteps to reach local waypoints. We develop a novel trajectory evaluation metric to minimize motion deviation and maximize information gain of the terrain elevation map. We evaluate the efficacy of our planning framework on Digit bipedal robot simulation in MuJoCo.iiiVideos of the simulated experiments in Mujoco can be found at https://youtu.be/27hOcBNKAvU?si=dp3GSGHndtQ7ZMaC
@@ II-C1 RBF kernel
The RBF kernel is a stationary kernel, commonly adopted in GP regression, in which the predictions are prone to be smooth and have the same degree of variability. The kernel is defined as
@@ II-C2 NN kernel
The NN kernel is nonstationary and models the covariance function of a neural network featuring a single hidden layer with infinitely many nodes and a sigmoid activation function [22]. The NN kernel is defined as
@@ II-C3 Attentive Kernel
The Attentive Kernel is nonstationary and adaptive to different variability in the output by applying a weighted sum of many base RBF kernels with different lengthscales. The kernel also utilizes the Instance Selection method, whereby each input location is assigned a membership vector to ensure that training input within the same neighborhood can break their correlation given abrupt changes in the training output activation function. The kernel is defined as
$\kappa(\bm{\xi}^{i},\bm{\xi}^{j})=\alpha\bar{z}^{T}\bar{z}^{\prime}+\sum_{m=1}^{M}\bar{w}_{m}\kappa_{m}(\bm{\xi}^{i},\bm{\xi}^{j})\bar{w}_{m}^{\prime}$
@@ IV-B Locomotion-dynamics-aware Local RRT*
$\begin{bmatrix}x_{{\rm apex}}^{\prime}&y_{{\rm apex}}^{\prime}\end{bmatrix}=$ (2)
$d_{{\rm safe}}\frac{\begin{bmatrix}x_{{\rm rand}}&y_{{\rm rand}}\end{bmatrix}-\begin{bmatrix}x_{\rm apex,1}&y_{\rm apex,1}\end{bmatrix}}{\left\|\begin{bmatrix}x_{{\rm rand}}&y_{{\rm rand}}\end{bmatrix}-\begin{bmatrix}x_{\rm apex,1}&y_{\rm apex,1}\end{bmatrix}\right\|_{2}},$ (7)
@@
$\begin{bmatrix}x_{{\rm apex}}^{\prime\prime}\\ y_{{\rm apex}}^{\prime\prime}\end{bmatrix}$ $=\begin{bmatrix}x_{{\rm apex,near}}\\ y_{{\rm apex,near}}\end{bmatrix}+d_{{\rm safe}}\begin{bmatrix}\cos(\theta_{{\rm near}}+\Delta\theta_{{\rm safe}})\\ \sin(\theta_{{\rm near}}+\Delta\theta_{{\rm safe}})\end{bmatrix},$ (14)
@@ V-A Locomotion-dynamics-aware Global RRT*
$lsb_{1}$ =$=\bigg{\{}\begin{bmatrix}x_{0}\\ y_{0}\end{bmatrix}+2d_{{\rm step}}\gamma\begin{bmatrix}\cos(\theta_{0}-\Delta\theta_{{\rm safe}})\\ \sin(\theta_{0}-\Delta\theta_{{\rm safe}})\end{bmatrix}\bigg{\}},\gamma\in[0,1],$ (19)
$lsb_{2}$ =$=\bigg{\{}\begin{bmatrix}x_{0}\\ y_{0}\end{bmatrix}+2d_{{\rm step}}\gamma\begin{bmatrix}\cos(\theta_{0}+\Delta\theta_{{\rm safe}})\\ \sin(\theta_{0}+\Delta\theta_{{\rm safe}})\end{bmatrix}\bigg{\}},\gamma\in[0,1],$ (24)
@@
$\begin{bmatrix}x^{\prime}&y^{\prime}\end{bmatrix}$ $=d_{{\rm step}}\frac{\begin{bmatrix}x_{{\rm rand}}&y_{{\rm rand}}\end{bmatrix}-\begin{bmatrix}x&y\end{bmatrix}}{\left\|\begin{bmatrix}x_{{\rm rand}}&y_{{\rm rand}}\end{bmatrix}-\begin{bmatrix}x&y\end{bmatrix}\right\|_{2}}$ (30)
@@ VI Results
We evaluate our framework on simulations of a Digit bipedal robot navigating three environments with varying terrain. Each environment is $20\times 20$ meters in size, with terrain elevation varying between 0 and 0.5 meters.
Figure 7: A sample run of our proposed framework using an Attentive-Kernel-based GP. In each graph, the red line represents the CoM trajectory; the yellow dots are the discrete footsteps; and the yellow star is the global goal location. (a) The ground truth elevation map with the heatmap. Regions with elevation above 0.15 meters are untraversable. (b) The final GP estimation of the terrain for the Attentive Kernel method. (c) The standard deviation of the terrain GP. Higher values indicate a higher uncertainty. Note that, as expected, the robot has a high confidence around its actual trajectory and low confidence far from where it has sampled. (d) The absolute error of the terrain GP (b) compared to the ground truth (a). The estimation shows high accuracy near the actual trajectory, and the robot does not need an accurate estimation of terrain far away from its trajectory in order to reach the goal. (e) A 3D view of the Digit robot navigation trajectory through the environment.
The terrain GP $\hat{z}$ model is initialized on prior information consisting of 100 evenly-spaced points across the terrain. Then, the GP model is updated online as new data is collected as the robot traverses the environment. The sensor is programmed to simulate the collection of 10 sample points within a 3-meter radius of the robot during each step. The model error GP $\Delta\hat{y}_{1}$ is trained offline as described in Section IV-A. Across all simulations, the average motion perturbation per step was 1.75e-2 meters, and the average prediction error of $\Delta\hat{y}_{1}$ per step was 2.06e-4 meters. Thus, $\Delta\hat{y}_{1}$ accurately evaluates the motion perturbation. Finally, for the LDA-L-RRT* local navigation planner, we evaluate three candidate trajectories for each local target.
@@
We also implement the simulation using full-order dynamics in MuJoCo, as depicted in Figure 1. We use a variation of the angular momentum LIP planner [24] to track the PSP plans as introduced in [20]. PSP hyperparameters (e.g., CoM velocities, and heading change) are used to design full-body joint trajectories through geometric inverse kinematics. A passivity-based controller [25] is used for full-body trajectory tracking. The supplemental video for this work shows Digit navigating through multiple environments using our framework.
@@
Avg. Error (Path) 3.05e-3 7.73e-4 2.02e-4
Avg. Error (Env.) 78.9 81.4 126
Avg. Time (sec) 231 167 310
Avg. Steps 196 188 161
Success Rate 90% 60% $<$53%
Finally, in Table I we benchmark the efficacy of the Attentive Kernel, RBF, and NN Kernel GPs in our case studies. For each of the three environments, we run experiments using each GP until we achieve three successes. Overall, the Attentive Kernel framework exhibits the highest success rate of reaching the global goal and minimizes the prediction error on the global terrain map (Env.), i.e., the entire environment.
The RBF kernel is the least expensive to implement computationally and achieves the lowest prediction error along the traversed trajectory (Path), i.e., where the robot actually stepped, but has a lower success rate than the Attentive Kernel. Finally, the NN Kernel requires the fewest number of steps on average to reach the global goal but it fails to find a successful path on one of the environments, thus making it the least reliable. The NN kernel, although non-stationary, correlates test points with both close-by and far-off points in the terrain, thus hindering performance in regions with rapidly varying terrain.

### v2

@@ Abstract
We study the problem of bipedal robot navigation in complex environments with uncertain and rough terrain. In particular, we consider a scenario in which the robot is expected to reach a desired goal location by traversing an environment with uncertain terrain elevation. Such terrain uncertainties induce not only untraversable regions but also robot motion perturbations. Thus, the problems of terrain mapping and locomotion stability are intertwined. We evaluate three different kernels, including two non-stationary kernels, for Gaussian process (GP) regression to learn the terrain elevation. We also learn the motion deviation resulting from both the terrain as well as the discrepancy between the reduced-order Prismatic Inverted Pendulum Model used for planning and the full-order locomotion dynamics. We propose a hierarchical locomotion-dynamics-aware sampling-based navigation planner. The global navigation planner plans a series of local waypoints to reach the desired goal locations while respecting locomotion stability constraints. Then, a local navigation planner is used to generate a sequence of dynamically feasible footsteps to reach local waypoints. We develop a novel trajectory evaluation metric to minimize motion deviation and maximize information gain of the terrain elevation map. We evaluate the efficacy of our planning framework on Digit bipedal robot simulation in MuJoCo. [^1] Videos of the simulated experiments in Mujoco can be found at https://youtu.be/27hOcBNKAvU?si=dp3GSGHndtQ7ZMaC
@@ II-C1 RBF kernel
The RBF kernel is a stationary kernel, commonly adopted in GP regression, in which the predictions are prone to be smooth and have the same degree of variability. The RBF kernel is defined as
@@ II-C2 NN kernel
The NN kernel is nonstationary and resembles a neural network featuring a single hidden layer with infinitely many depth nodes and a sigmoid activation function [22]. This kernel models spatial correlation between data points which depends on the distance from the data origin until reaching a saturation region. The datapoints within the saturation region also correlate among themselves, allowing this kernel to adapt to rapid variations in the output values.
The work [14] leverages the NN kernel for large-scale terrain elevation reconstruction, utilizing a local approximation method applied to the kernel. The approach involves implementing a KD-Tree to efficiently search for a specific number of nearest neighbors at each query point, facilitating the prediction of terrain elevation. This work demonstrates the suitability of the NN kernel for predicting terrain characterized by high variability in elevation and widely dispersed training data. This is due to the kernel’s treatment of each query point as a data origin, correlating the output prediction solely with the training data nearest to that origin before reaching the saturation region. Within the saturation region, the kernel independently correlates other data points.
Despite the efficacy demonstrated by the terrain mapping method introduced in [14], incorporating such a local approximation technique into our framework entails significant computational overhead. The issue arises from the need to identify unique set of nearest neighbors for each query point across the terrain and individually compute predictions for each query point, thus significantly increasing the computation time. Consequently, we propose a K-means clustering strategy to group closer input locations together and separate them into k clusters. This approach ensures that input locations within the same cluster share identical sets of nearest neighboring training data, thereby mitigating the computational load. Detailed block diagram of the local approximation method is shown in Fig 3. In this study, we opt to partition our input space, comprising 2,500 locations (as detailed in VI), into $k=200$ clusters to expedite terrain elevation prediction.
The NN kernel is defined as
@@ II-C3 Attentive Kernel
The Attentive Kernel is nonstationary and adaptive to different variability in the output by applying a weighted sum of many base RBF kernels with different lengthscales. The kernel also utilizes the Instance Selection method, whereby each input location is assigned a membership vector to ensure that the training input within the same neighborhood can break their correlation given abrupt changes in the training output [13]. The incorporation of weighted summation and instance selection allows the Attentive kernel to achieve robust predictive performance in the face of terrain elevation variability, while also enhancing its capability for uncertainty quantification. Both weights and membership vectors are learned from the training dataset using a two-hidden-layer neural network with TanH activation function. The kernel is defined as
$\kappa(\bm{\xi}^{i},\bm{\xi}^{j})=\alpha\bar{z}^{T}\bar{z}^{\prime}\sum_{m=1}^{M}\bar{w}_{m}\kappa_{m}(\bm{\xi}^{i},\bm{\xi}^{j})\bar{w}_{m}^{\prime}$
@@
Figure 3: Overall block diagram of the local approximation method for GP prediction with NN kernel. Whenever the global elevation map undergoes re-estimation, all input locations in the terrain are split into k clusters via K-mean clustering algorithm. Each cluster consists of a centerpoint $c_{i}$ and the subset input locations $\{\bm{\xi}\}_{i}$. Iterating through each of the k clusters, the centerpoint is passed to the KD-Tree method to find n nearest neighboring training datapoints denoted as $D_{nearest}^{i}=\{(\bm{\xi}^{j},y^{j})\}^{i}_{j\in\{1,...,n\}}$. Given n closest training datapoints to the centerpoint (i.e., $D_{nearest}^{i}$) and the set of test points in that cluster to be estimated (i.e., $\{\bm{\xi}\}_{i}$), we then have all of the information to make GP predictions of the terrain heights at all input locations within the $i^{th}$ cluster region. With this local approximation method, GP prediction with NN kernel is only computed k times, thereby reducing the computational load as compared to making a separate prediction for all input locations.
@@ IV-B Locomotion-dynamics-aware Local RRT*
$\begin{bmatrix}x_{{\rm apex}}^{\prime}&y_{{\rm apex}}^{\prime}\end{bmatrix}=$
$d_{{\rm safe}}\frac{\begin{bmatrix}x_{{\rm rand}}&y_{{\rm rand}}\end{bmatrix}-\begin{bmatrix}x_{\rm apex,1}&y_{\rm apex,1}\end{bmatrix}}{\left\|\begin{bmatrix}x_{{\rm rand}}&y_{{\rm rand}}\end{bmatrix}-\begin{bmatrix}x_{\rm apex,1}&y_{\rm apex,1}\end{bmatrix}\right\|_{2}},$
@@
$\begin{bmatrix}x_{{\rm apex}}^{\prime\prime}\\ y_{{\rm apex}}^{\prime\prime}\end{bmatrix}$ $=\begin{bmatrix}x_{{\rm apex,near}}\\ y_{{\rm apex,near}}\end{bmatrix}+d_{{\rm safe}}\begin{bmatrix}\cos(\theta_{{\rm near}}+\Delta\theta_{{\rm safe}})\\ \sin(\theta_{{\rm near}}+\Delta\theta_{{\rm safe}})\end{bmatrix},$
@@ V-A Locomotion-dynamics-aware Global RRT*
$lsb_{1}$ =$=\bigg{\{}\begin{bmatrix}x_{0}\\ y_{0}\end{bmatrix}+2d_{{\rm step}}\gamma\begin{bmatrix}\cos(\theta_{0}-\Delta\theta_{{\rm safe}})\\ \sin(\theta_{0}-\Delta\theta_{{\rm safe}})\end{bmatrix}\bigg{\}},\gamma\in[0,1],$
$lsb_{2}$ =$=\bigg{\{}\begin{bmatrix}x_{0}\\ y_{0}\end{bmatrix}+2d_{{\rm step}}\gamma\begin{bmatrix}\cos(\theta_{0}+\Delta\theta_{{\rm safe}})\\ \sin(\theta_{0}+\Delta\theta_{{\rm safe}})\end{bmatrix}\bigg{\}},\gamma\in[0,1],$
@@
$\begin{bmatrix}x^{\prime}&y^{\prime}\end{bmatrix}$ $=d_{{\rm step}}\frac{\begin{bmatrix}x_{{\rm rand}}&y_{{\rm rand}}\end{bmatrix}-\begin{bmatrix}x&y\end{bmatrix}}{\left\|\begin{bmatrix}x_{{\rm rand}}&y_{{\rm rand}}\end{bmatrix}-\begin{bmatrix}x&y\end{bmatrix}\right\|_{2}}$
@@ VI Results
We evaluate our framework on simulations of a Digit bipedal robot navigating three environments with varying terrain (i.e., N37W112, N24W102, N17E102) as shown in Figure 9. Each of the three terrains exhibits distinct levels of nonstationarity. The N37W112 terrain is characterized by numerous standalone hills that render it impassable for the bipedal robot, presenting a high degree of nonstationarity marked by many abrupt changes in elevation. In contrast, the N24W102 terrain features fewer significant changes in elevation, displaying a more gradual form of nonstationarity. Finally, the N17E102 terrain exhibits the least nonstationarity, consisting of only one elongated elevation across the landscape. Each environment is $20\times 20$ meters in size, with terrain elevation varying between 0 and 0.5 meters.
Figure 8: A sample run of our proposed framework using an Attentive-Kernel-based GP on the terrain N37W112. In each graph, the red line represents the CoM trajectory; the yellow dots are the discrete footsteps; and the yellow star is the global goal location. (a) The ground truth elevation map with the heatmap. Regions with elevation above 0.15 meters are untraversable. (b) The final GP estimation of the terrain for the Attentive Kernel method. (c) The standard deviation of the terrain GP. Higher values indicate a higher uncertainty. Note that, as expected, the robot has a high confidence around its actual trajectory and low confidence far from where it has sampled. (d) The absolute error of the terrain GP (b) compared to the ground truth (a). The estimation shows high accuracy near the actual trajectory, and the robot does not need an accurate estimation of terrain far away from its trajectory in order to reach the goal. (e) A 3D view of the Digit robot navigation trajectory through the environment.
(a) N37W112
(b) N24W102
(c) N17E102
Figure 9: Three environments used in the navigational task case study with sample final trajectory (start from white to red).
The terrain GP $\hat{z}$ model is initialized on prior information consisting of 500 randomly-selected points across the terrain (out of the total of 2500 points). Then, the GP model is updated online as new data is collected as the robot traverses the environment. The sensor is programmed to simulate the collection of 10 sample points within a 3-meter radius of the robot during each step. The model error GP $\Delta\hat{y}_{1}$ is trained offline as described in Section IV-A. Across all simulations, the average motion perturbation per step was 1.75e-2 meters, and the average prediction error of $\Delta\hat{y}_{1}$ per step was 2.06e-4 meters. Thus, $\Delta\hat{y}_{1}$ accurately evaluates the motion perturbation. Finally, for the LDA-L-RRT* local navigation planner, we evaluate three candidate trajectories for each local target.
@@
We also implement the simulation using full-order dynamics in MuJoCo, as depicted in Figure 1. We use a variation of the angular momentum LIP planner [24] to track the PSP plans as introduced in [20]. PSP hyperparameters (e.g., CoM velocities, and heading change) are used to design full-body joint trajectories through geometric inverse kinematics. A passivity-based controller [25] is used for full-body trajectory tracking. The supplemental video for this work shows Digit navigating through multiple environments using our framework.11footnotemark: 1
@@
Avg. Error (Path) 2.04e-4 1.10e-3 2.83e-4
Avg. Std. Dev. (Path) 2.58e-2 2.67e-2 5.19e-2
Avg. Error (Env.) 52.60 49.61 49.02
Avg. Std. Dev. (Env.) 3.13e-2 3.13e-2 5.48e-2
Avg. Time [sec] 303 313.48 352.42
Avg. Steps 638 617 667
Retrain every # steps 20 14 19.33
In Table I, we benchmark the efficacy of the Attentive Kernel, RBF, and NN Kernel GPs in our case studies. For each of the three environments, we run experiments using each GP until we achieve three successful runs. This is achieved by incrementally increasing the retraining frequency of the GP (measured in number of walking steps taken by the robot). This iterative approach ensures that the GP model can provide accurate predictions of the local terrain elevation, thus enabling the bipedal robot to navigate locally without encountering impassable terrain heights. Overall, the Attentive kernel demonstrates the lowest prediction error along the traversed trajectory (Path), i.e., the path where the robot actually traversed, while the Neural Network kernel (with KD-tree and K-mean) excels in minimizing the prediction error across the global terrain map (Env.), i.e., the entire environment. Furthermore, the Attentive kernel showcases notable proficiency in uncertainty quantification, evident in both the local traversed trajectory and the global terrain map, with the lowest average standard deviation for each prediction. Both the Attentive Kernel and the Neural Network kernel exhibit comparable retraining frequencies, allowing for a greater number of walking steps before updating the GP model. However, the Attentive kernel emerges as the more computationally efficient option, characterized by a lower average total trajectory time. Lastly, although the RBF kernel facilitates the fewest total walking steps to reach the goal, it requires more frequent retraining of the GP model to achieve three successful runs.
Figure 10: Benchmarking of three considered GP kernels: (a) Average mean prediction error across the entire environment [(Ground Truth) - (GP estimated mean)] (b) Average standard deviation across the entire environment (c) Average GP prediction times (d) Average GP retraining times
In Figure 10, we conducted additional experiments on terrain N37W112, averaging over 10 successful trials for each kernel. First, our findings reveal that the Neural Network kernel demonstrates the lowest mean prediction error for estimating the global terrain map, albeit with the highest uncertainty (average standard deviation) among the three kernels. Conversely, the Attentive Kernel exhibits superior performance in uncertainty quantification with the lowest average standard deviation throughout all local planning epochs, consistent with our previous benchmarking results, and displays enhanced prediction accuracy compared to the baseline RBF kernel.
With respect to computational complexity, the Neural Network kernel initially shows higher prediction times due to its utilization of a local approximation method, as discussed in II-C. This method necessitates separate predictions for each cluster of query points and results in an extended duration for multiple predictions. However, the local approximation methods enable the Neural Network kernel to remain robust against the accumulation of newly-collected training data, as it consistently requires the same number of training datapoints for prediction. Consequently, the prediction time remains relatively consistent regardless of the total number of training datapoints amassed. In contrast, both the Attentive kernel and the RBF kernel initially demonstrate rapid prediction computation but increase in prediction time as the number of training samples grows. This scaling arises from both kernels’ utilization of all training data to predict terrain elevations, leading to their prediction times eventually exceeding those of the Neural Network kernel around the $9^{\rm th}$ local planning epoch. Lastly, the training time for the Neural Network kernel and the RBF kernel is comparable, whereas the Attentive kernel necessitates additional training time for optimizing the hyperparameters for weighted summation of its various base kernels and for optimizing the membership vector assigned to each input location.
In summary, the Attentive kernel excels in uncertainty quantification and locally-traversed terrain elevation prediction, while the Neural Network kernel demonstrates better performance in predicting the entire environment despite exhibiting higher uncertainty. For long-horizon trajectory planning (or trajectory longer than our case study), the Neural Network kernel combined with a local approximation method is recommended for terrain GP modeling due to its consistent computational time for terrain estimation at every epoch. While the Attentive kernel can also be employed for long-horizon trajectories, the implementation of data allocation methods is advised to alleviate the computational complexity during the prediction phase. The same local approximation methods may be applied to the Attentive kernel, although the prediction accuracy and uncertainty quantification performance can potentially be compromised. Overall, both the Neural Network kernel and the Attentive kernel demonstrate superior performance compared to the baseline RBF kernel, attributed to their adaptability to the non-stationary nature of terrain.